If 2 cars were 300 miles apart at 12 noon, and drove until they met each other at 3 pm, find the speeds of both cars. 1 car was driving at a speed of 10 mph faster then the other.
Answer by ikleyn(52812)

.

Let x be the rate of the slower car, in miles per hour;

then the rate of the faster car is (x+10) mph.


The total distance equation is then


    3x + 3*(x+10) = 300  miles   (the "3" here is "3 hours from 12 noon to 3 pm).


From the equation,


    3x + 3x + 30 = 300

    6x           = 300 - 30 = 270

     x                      = 270/6 = 45.


ANSWER.  The slower car rate was  45 mph;  the faster car  45+10 = 55 mph.

Solved.

----------------

See introductory lessons on Travel and Distance problems
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.